OpenStudy (gatsbyyy):

The perimeter of a triangle is 147 inches. If one side of the triangle is five more than the shortest side, and the longest side is 16 more than the shortest side, find the lengths of the three sides? Shortest side= Middle-valued side= Longest side=

OpenStudy (gatsbyyy):

I tried making an equation but the numbers came out weird Medium-valued side: 147=x+5 I did the equation but then I got 142 which doesn't make sense because I don't have the sides for the other equations.

jimthompson5910 (jim_thompson5910):

x = shortest side y = middle-length side z = longest side x < y < z `The perimeter of a triangle is 147 inches` x+y+z = 147 `If one side of the triangle is five more than the shortest side` y = x+5 `the longest side is 16 more than the shortest side,` z = x+16

jimthompson5910 (jim_thompson5910):

x+y+z = 147 x+(x+5)+z = 147 ... plug in y = x+5 x+x+5+z = 147 x+x+5+(x+16) = 147 ... plug in z = x+16 x+x+5+x+16 = 147 3x+21 = 147

jimthompson5910 (jim_thompson5910):

I'll let you finish up

OpenStudy (gatsbyyy):

I got 42 for x

jimthompson5910 (jim_thompson5910):

x = 42 is correct. Now use this to find y and z